# FFHFlow: A Flow-based Variational Approach for Learning Diverse Dexterous Grasps with Shape-Aware Introspection

### v2

@@ V-F Evaluation in the Real World
Unconfined space
This setup is similar to the simulation, where 8 objects are grasped 10 times each. For Q3, the results in Table IV show that FFHFlow-lvm  is able to generalize to the real hardware and real-world data much better in terms of average success rate, namely $15.0\%$ increase to the FFHNet baseline. Some example grasps are visualized in 8.
Confined space
We selected a confined space, i.e., a two-tier shelf to further mimic the realistic scenarios (d) in Figure 4. In Table IV, a large performance gain is achieved by FFHFlow-lvm when compared to cVAE. The main failure of cVAE is due to its biased top grasps colliding with the environment (shelf). In such a confined space, a diverse grasp synthesizer based on FFHFlow-lvm is more effective than the mode-seeking cVAE.
Cluttered Scenes
We conduct grasping experiments in cluttered scenes as (e) and (f) shown in Figure 4. The generated grasps are sent to FCL to filter out grasps that are in collision with the whole point cloud. In Table IV, FFHFlow-lvm achieves a higher success rate (SR) and also clearance rate (CR) compared to baseline.
V-G Ablation study
V-G1 Flow-based Approaches
We evaluate different model variations in Table III. The improved performance of FFHFlow-lvm over FFHFlow-cnf can be attributed to its more informative latent features leveraging DLVMs. To isolate the effect of model size on performance, we introduce FFHFlow-lvm-light, a streamlined version of FFHFlow-lvm in which both Prior Flow and Grasp Flow consist of 4 layers instead of 8. This FFHFlow-lvm-light has a model size and runtime comparable to FFHFlow-cnf, yet it still achieves superior performance, with a $1.3\%$ improvement on similar objects and $12.2\%$ on novel objects.
TABLE V: $\epsilon$ in Introspective Grasp Evaluation
Additive Coefficients ($\epsilon$) $0.0$ $0.01$ $0.1$ $0.5$ $1.0$ Similar $90.5\%$ 94.6% $90.6\%$ $78.6\%$ $63.0\%$ Novel $50.9\%$ 52.7% $50.9\%$ $34.3\%$ $25.3\%$
V-G2 Introspective Grasp Evaluation
We conduct an ablation study presented in Table II to understand the trade-off between increasing grasp quality (grasp evaluator) and lowering view-level shape uncertainty (Grasp Flow), namely the optimal value of the additive coefficient ($\epsilon$). By increasing the impact of lowering $\epsilon$, we can see the performance first increases and drops. The optimal value is $0.01$, indicating the major contribution to grasp success from the grasp evaluator.
VI Conclusion
We introduce a novel flow-based variational approach, FFHFlow-lvm for generative grasp synthesis with better quality and diversity. This is achieved by mitigating the insufficiently informative latent features when applying cNF directly and overcoming problems in cVAE-based approaches, i.e., mode-collapse and the mis-specified prior. Moreover, the model is equipped with shape-aware introspection quantified by the exact flow likelihoods, which further facilitates a novel hybrid scheme for uncertainty-aware grasp evaluation. Comprehensive experiments in the simulation and real world demonstrate strong performance and efficiency.

### v3

@@ 5.4 Evaluation in the Real World
Unconfined Space: This setup is similar to the simulation, where 8 objects are grasped 10 times each. For Q4, Table 2 shows that FFHFlow-lvm  is able to generalize to the real robot with a smaller gap. Confined Space: We selected a confined space, i.e., a two-tier shelf to further mimic the realistic scenarios (d) in Figure 4(a). In Table 2, a large performance gain is achieved by FFHFlow-lvm when compared to cVAE. The main failure of cVAE is due to its biased top grasps colliding with the environment (shelf). In such a confined space, a diverse grasp synthesizer is more effective than the mode-seeking cVAE. Cluttered Scenes: We conduct grasping experiments in cluttered scenes as (e) and (f) shown in Figure 4(a). The metrics Clearance Rate (CR) denotes the probability of robots clear the scene. The predicted grasps with collision are filtered out based on FCL. FFHFlow-lvm achieves a higher success rate (SR) and also clearance rate (CR). More details in the appendix.
6 Conclusion
We introduce a novel flow-based variational approach, FFHFlow-lvm for generative grasp synthesis with better quality and diversity. This is achieved by mitigating the insufficiently informative latent features when applying cNF directly and overcoming problems in cVAE-based approaches, i.e., , mode-collapse and the mis-specified prior, as well as inefficiency issues from diffusion-based approaches. Moreover, the model is equipped with shape-aware introspection quantified by the exact flow likelihoods, which further facilitates a novel hybrid scheme for uncertainty-aware grasp evaluation. Comprehensive experiments in the simulation and real world demonstrate strong performance and efficiency.
7 Limitations
Though our proposed idea exhibits numerous strengths, its limitations remain to solve for further improvement in the future. We highlight the most pronounced limitations to facilitate future research. (1) Trade-off between run-time and performance: reducing the flow size can improve the run-time but at expense of a slight performance loss. Investigating how to strike a balance is relevant for broader robotic applications. (2) The lack of abilities for efficient adaptation towards objects that differ significantly from those in the training dataset, which is hard to avoid for robots deployed in the wild [52, 59, 60]. We can further scale up the synthetic dataset, leveraging GPU-powered simulation such as Isaac Gym/Sim to increase generalization capability. (3) Sim-to-Real gap: Though the objects in simulation and real world are different in evaluation, a success rate drop of $17.1\%$ encourages us to further investigate how to eliminate the sim-to-real gap, through various approaches such as camera noise modeling, physical parameters randomization and better simulators.